Give one example of solving a problem using mgf of normal distribution
1
Expert's answer
2021-08-20T12:08:58-0400

A continuous random variable X is said to have a Normal Distribution if its PDF is:

"f_X(x)=\frac{1}{σ \sqrt{2π}}e^{−\frac{1}{2}(\frac{x−μ}{σ})^2},x∈R\\"

So if X∼N(μ,σ2) then it’s MGF is:

"M_x(t)\\ 

=E(e^{tx})\\

=∫^∞_{−∞}e^{tx}\frac{1}{σ \sqrt{2π}}e^{−\frac{1}{2}}(\frac{x−μ}{σ})^2dx\\

=∫^∞_{−∞}e^t(μ+σz)\frac{1}{\sqrt{2π}}e−\frac{1}{2}z^2dz, where \frac{x−μ}{σ}=z\\

=(e^{tμ}+\frac{1}{2}t^2σ^2)∫^∞_{−∞}\frac{1}{\sqrt{2π}}e^−\frac{1}{2}(z−tσ)^2dz,"

Note that the portion under integration sign represents PDF of N(tσ,1) distribution; so eventually the value of the integration will be 1

"e^{tμ}+\frac{1}{2}t^2σ^2, t ∈R"

So, "M_x(t)= e^{tμ}+\frac{1}{2}t^2σ^2, t ∈R"